1. what is the [H+] if the [OH-]=7.1x10^-8?
2. what is the pH of a solution with [OH-]=2.1x10^-4?
3. what is the [H+] if the pH=3.17?
4. what is the [OH-] if the pH=7.20?
Answer #1

1) we know poh = -log[OH-] substitiute value we get poh = -log(7.1x10-8) = 7.14

ph = 14 - poh = 14- 7.14 = 6.86  

we know formula [H+] = 10-ph = 10-6.86 =1.38 x 10-7

ans [H+] = 1.38 x 10-7 M

2) we know formula poh = -log[OH-] = -log(2.1x10-4) = 3.6

pH = 14 - pOH = 14-3.6 = 10.4

ans PH = 10.6  

3) we know formula [H+] = 10-PH substitute value [H+] = 10-3.17 = 6.76x10-4

ans [H+] = 6.76x10-4 M

4) we know formula POH = 14 - PH = 14-7.2 = 6.8

we know formula [OH-] = 10-POH substitute value [OH] = 10-6.8 = 1.58x10-7

ans [OH-] = 1.58x10-7M